40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 9 12:35:47 2017, gautam, Update, ALS, Fiber ALS IMG_3942.JPG
    Reply  Mon Aug 14 16:24:09 2017, gautam, Update, ALS, Fiber ALS IMG_7471.JPGIMG_7472.JPGIMG_7473.JPGIMG_7474.JPG
       Reply  Wed Aug 16 20:24:23 2017, gautam, Update, ALS, Fiber ALS IMG_7476.JPGIMG_7477.JPGIMG_7478.JPGIMG_7479.JPG
          Reply  Wed Aug 23 17:22:36 2017, gautam, Update, ALS, Fiber ALS - reinstalled 
             Reply  Fri Aug 25 15:54:14 2017, gautam, Update, ALS, Fiber ALS noise measurement IMG_7497.JPGFOL_schematic.pdf20170825_IR_ALS.pdf
                Reply  Fri Aug 25 17:11:07 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Sep 26 19:10:13 2017, gautam, Update, ALS, Fiber ALS setup neatened IMG_7605.JPG
                Reply  Sun Aug 27 11:57:31 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Aug 29 02:08:39 2017, gautam, Update, ALS, Fiber ALS noise measurement 
                      Reply  Fri Sep 1 19:15:40 2017, gautam, Update, ALS, Fiber ALS noise measurement IR_ALS_noise.pdf
Message ID: 13288     Entry time: Fri Sep 1 19:15:40 2017     In reply to: 13266
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber ALS noise measurement 

Summary:

I did some work today to see if I could use the IR beat for ALS control. Initial tests were encouraging.

I will now embark on the noise budgeting.

Details:

  • For this test, I used the X arm
  • I hooked up the X-arm + PSL IR beat to the X-arm DFD channel, and used the Y-arm DFD channels to simultaneously monitor the X-arm green beat.
  • I then transitioned to ALS control and used POX as an out-of-loop sensor for the ALS noise.
  • Attachment #1 shows a comparison of the measurements. In red is the IR beat, while the green traces are from the test EricQ and I did a couple of nights ago using the green beat.
  • I also wanted to do some arm cavity scans with the arm under ALS control with the IR beat - but was unsucessful. The motivation was to fix the ALS model counts->Hz calibration factors.
  • I did however manage to do a 10 FSR scan using the green beatnote - however, towards the end of this scan, the green beat frequency (read off the control room analyzer) was ~140MHz, which I believe is outside (or at least on the edge) of the bandwidth of the Green BBPDs. The fiber coupled IR beat photodiodes have a much larger (1GHz) spec'd bandwidth.

I am leaving the green beat electronics on the PSL table in the switched state for further testing...

 

Attachment 1: IR_ALS_noise.pdf  39 kB  Uploaded Sat Sep 2 16:42:21 2017  | Hide | Hide all
IR_ALS_noise.pdf
ELOG V3.1.3-